# DeSTA2.5-Audio: Toward General-Purpose Large Audio Language Model with Self-Generated Cross-Modal Alignment

### v1

@@ VII Comparison Studies
As shown in Table III, we evaluate the proposed self-generation approach under controlled experimental setups using identical audio-description pairs from a 500K subset of DeSTA-AQA5M. The results are analyzed across four settings to provide a comprehensive understanding. The self-generation setting represents the proposed configuration, where the same backbone LLM is used both to construct the training target and to serve as the LALM backbone. In this case, there is no distribution mismatch between the model and the training data. In the cross-model setting, different LLMs are used to generate the training targets. This setup simulates a distribution mismatch, such as when human annotators or alternative LLMs are employed to create the training corpus. This setup is widely adopted in previous LALM studies [8, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Building upon these two setups, we explore configurations with LoRA adapters integrated into the backbone LLM. This allows us to examine whether introducing additional trainable parameters helps mitigate performance degradation due to data mismatch or enhances generalization. Lastly, to examine the effect of training duration on model performance, we adopt a 5-epoch setting in contrast to our default 10-epoch setup. This allows us to investigate how increased exposure to the training data influences model performance. To quantify the degree of distribution mismatch, we report the perplexity (PPL) of the training targets as computed by the corresponding backbone LLM. A lower perplexity indicates the model is more familiar with the training target, suggesting less distributional discrepancy.
VII-A Comparison between Self-generated and Cross-model Settings
@@
As indicated in Table III, the self-generated training data exhibit consistently lower perplexity, suggesting that the generated responses align well with the distribution of the backbone LLM. When comparing Llama3.1 (A1) and Qwen2.5 (A2), Qwen2.5 consistently outperforms Llama3.1 across all benchmarks. This performance gap may be attributed to Qwen2.5’s stronger text generation capabilities. While Qwen2.5’s performance in basic content comprehension tasks is relatively on par with Llama3.1, it demonstrates better performance in other areas, such as speaker category in Dynamic-SUPERB Phase-1, and environmental sound and music understanding in MMAU. Previous evaluations on text-based benchmarks have also shown that Qwen2.5 exhibits superior reasoning and mathematical abilities compared to Llama3.1 [121]. However, there is currently no conclusive evidence indicating a corresponding advantage in auditory perception, which warrants further investigation. Nevertheless, under identical training conditions, our experimental results suggest that Qwen2.5 serves as a more effective backbone LLM than Llama3.1. These findings also indicate that our training framework generalizes well across different LLMs.
Prompt diversity also plays a significant role in model performance, particularly in the comparison between A1 and A3. In A3, we adopt the self-generation setup using a single descriptive prompt (1-p), as described in our previous work [9], which already demonstrates strong zero-shot generalization. By simply increasing prompt diversity, as done in A1, we further enrich the training target and enhance the overall effectiveness of the training method. Notably, these results are achieved without the need for any task-specific instruction pairs. This highlights the strength of the self-generation design. Even when data construction relies solely on randomly sampled prompts, the model can still achieve zero-shot generation by leveraging the inherent capabilities of the LLM.
When comparing self-generation and cross-model settings, training targets in the cross-model setting result in higher perplexity, suggesting that the backbone LLM is less familiar with the data distribution. For example, while training Qwen2.5 on Qwen2.5-generated data (A2) produces strong results, training Llama3.1 on Qwen2.5-generated data (B1) leads to model degeneration [122], with outputs containing repetitive or nonsensical tokens. Similarly, training Llama3.1 on data generated by Gemma3-12B (B2) fails to match the performance observed in the self-generation setting (A1). These results support our distribution mismatch hypothesis and emphasize the importance of using a self-generated configuration, even when the annotator LLM is more capable. We also explore using Llama3.1-70B to generate training data (B3), representing a more powerful model from the same family. In this case, the lower perplexity (2.20) suggests a closer alignment between the training data and Llama3.1’s distribution. However, compared to A1, B3 achieves better performance on Dynamic-SUPERB and SAKURA, but underperforms on MMAU and Speech-IFEval. This indicates that using a stronger model does not necessarily lead to consistent improvements across all tasks.
@@ VII-B With LoRA Adapter
In the LoRA adapter setting, we introduce trainable parameters to the backbone LLM, which is expected to increase model capacity and help mitigate the distribution mismatch problem. In the self-generation setup (C1), where the dataset is well-aligned with the backbone LLM, we find that adding LoRA layers yields similar or only slightly better performance. This indicates that, under self-generation settings, incorporating LoRA adapters does not provide significant advantages. In other words, fine-tuning a lightweight modality adapter is sufficient for cross-modal alignment when using our proposed training framework, where the model focuses on learning auditory concepts without being hindered by stylistic or distributional mismatches. Interestingly, when training with Qwen2.5-generated data (C2), the performances in audio processing benchmarks, such as Dynamic-SUPERB, MMAU, and SAKURA-Single, are comparable to the self-generation setup (A2). However, they experience a significant degradation in SAKURA-Multi and Speech-IFEval, which requires additional text knowledge and instruction-following ability. This difference indicates that while adding a LoRA adapter can help mitigate the distribution mismatch and perform well on in-domain tasks, it may still degrade the model’s general capabilities when evaluated on benchmarks that require knowledge from LLM pretraining. This reveals a critical design limitation in current LALM training strategies. Models such as LTU-AS and SALMONN attempt to address the catastrophic forgetting problem by introducing LoRA adapter layers to the LLM. However, our experimental results suggest that reducing the discrepancy between training data and model distribution is a more critical factor for preserving generalization ability than architectural modifications alone.
VII-C With 5 epochs
With the 5-epoch setting, we investigate the impact of training duration on model performance. Consistent with discussion in Section VII-A, the 5-epoch results (D1 and D2) show that prompt diversity not only improves effectiveness but also enhances training efficiency. Despite being trained for only half the number of epochs, the models achieve performance comparable to their 10-epoch counterparts (A1). Notably, while D2 continues to improve with extended training (as in A3), convergence is slower and the final performance remains worse, indicating that diverse training target is also essential for achieving better alignment. In contrast, although D3 achieves non-trivial performance with only 5 epochs of training, B1 demonstrates that extended training under distribution mismatch leads to model degeneration. These findings underscore our key motivation: effective cross-modal alignment requires repeated exposure to align audio representations across epochs. When the training data is matched with the backbone model, performance improves steadily without degrading the model’s inherent language capabilities. In contrast, learning from mismatched data places a heavier burden on the model, ultimately resulting in suboptimal performance and forgetting of its pretrained linguistic abilities.
@@ VIII Conclusion
In this work, we introduced DeSTA2.5-Audio, a general-purpose LALM built on the self-generated cross-modal alignment framework, DeSTA. By allowing the backbone LLM to generate its own training targets, DeSTA maintains distributional consistency and mitigates the catastrophic forgetting problem commonly observed in prior works. Under this framework, we successfully develop a robust and generalizable LALM without relying on task-specific instruction data. We constructed DeSTA-AQA5M, a large-scale, task-agnostic dataset comprising 5 million audio-text pairs from 50 diverse audio datasets. Trained solely on this corpus, DeSTA2.5-Audio achieves state-of-the-art or competitive results across benchmarks such as Dynamic-SUPERB (Phase-1 and Phase-2), MMAU, SAKURA, Speech-IFEval, and VoiceBench. Our findings highlight the importance of training data quality over quantity in LALM development. Compared to previous studies, our self-generation strategy offers a scalable and robust solution with better generalization.
While our current design leverages text description as an intermediate bridge, we acknowledge that not all acoustic nuances can be effectively captured through textual representations. In future work, we will continue to explore methods that enable LALMs to better capture subtle, non-text-expressible audio features.

### v2

@@ VII Comparison Studies
As shown in Table V, we evaluate the self-generation approach under controlled experimental setups using identical audio-description pairs from a 500K subset of DeSTA-AQA5M. We compare the self-generation setting (backbone-generated targets) against external-model settings using targets from other LLMs. Notably, the external-model setup simulates a distribution mismatch, such as when human annotators or alternative LLMs are employed to create the training corpus, which is widely adopted in previous LALM studies [72, 35, 36, 92, 34, 47, 61, 18, 17, 1, 50, 33]. We also analyze the performance evolution in Figure 7 and cross-modal alignment quality in Table VI. Building upon these two setups, we explore configurations with LoRA adapters integrated into the backbone LLM. To quantify the degree of distribution mismatch, we report the perplexity (PPL) of the training targets as computed by the corresponding backbone LLM. A lower perplexity indicates the model is more likely to produce the training target sequence, suggesting less distributional discrepancy in the dataset.
Figure 7: Epoch-wise evolution of different data construction settings on instruction-based SAKURA-Single, which requires combined audio perception and instruction-following abilities.
VII-A Comparison between Self-generation and External-model
@@
As indicated in Table V and Figure 7, the self-generated setup exhibits consistently superior performance and training stability. Since the training targets are consistent with the backbone’s native output distribution, the optimization process avoids conflicting stylistic adaptation, focusing exclusively on cross-modal alignment. When comparing Llama3.1 (A1) and Qwen2.5 (A2), Qwen2.5 consistently outperforms Llama3.1 across all benchmarks. This performance gap may be attributed to Qwen2.5’s stronger text generation capabilities. While Qwen2.5’s performance in basic content comprehension tasks is relatively on par with Llama3.1, it demonstrates better performance in other areas, such as speaker category in Dynamic-SUPERB Phase-1, and environmental sounds and music understanding in MMAU. Previous evaluations on text-based benchmarks have also shown that Qwen2.5 exhibits superior reasoning and mathematical abilities compared to Llama3.1 [29]. However, there is currently no conclusive evidence indicating a corresponding advantage in auditory perception, which warrants further investigation. Nevertheless, under identical training conditions, our experimental results suggest that Qwen2.5 serves as a more effective backbone LLM than Llama3.1. These findings also indicate that our training framework generalizes well across different LLMs.
In contrast, external-model settings introduce distributional shifts that severely disrupt training. Figure 7 illustrates the severe consequences of this conflict. In the B1 setting (Llama3.1 trained on Qwen2.5-generated data), we observe catastrophic model degeneration [89] after the fourth epoch, manifesting as repetitive or nonsensical outputs. Despite Qwen2.5-generated data proving effective in self-generated experiments (A2), this failure suggests that the discrepancy destabilizes the training process. Similarly, while the B2 setting (Llama3.1 trained on Gemma3-generated data) avoids model collapse, the learning curve plateaus prematurely, indicating that the distributional discrepancy still hinders the model’s ability to refine auditory perception.
TABLE VI: Alignment quality analysis on SAKURA-Single comparing instruction-based QA (Inst.) and similarity-based classification (Sim.). L, Q and G denote Llama, Qwen and Gemma, respectively.
ID Gender Emotion Animal
Inst. Sim. Inst. Sim. Inst. Sim.
Random 50.0 20.0 11.1
A1 (L-L) 91.0 93.0 57.2 55.6 57.4 65.4
A2 (Q-Q) 89.8 86.8 62.2 54.8 77.8 77.0
B1 (L-Q) 28.0 58.6 5.8 20.4 30.4 15.2
B2 (L-G) 62.4 83.4 42.0 34.4 24.2 26.2
To determine whether these failures arise from poor alignment or generation issues, Table VI presents an analysis based on instruction-following and similarity-based metrics. The self-generated settings (A1, A2) achieve high scores in both metrics, demonstrating successful cross-modal alignment that maps audio features to semantic concepts within the representation space. Consequently, the model effectively leverages its intrinsic instruction-following capabilities to jointly process audio inputs and text instructions, facilitating robust instruction-based evaluation. Conversely, the external-model settings (B1, B2) exhibit significant degradation in both areas. Given that the training target is the sole variable, these results reveal that distributional discrepancy not only undermines training stability but also impedes the effective integration of audio information, directly leading to the poor downstream performance observed across all metrics in Table V.
Finally, comparing A1 and A3 (Table  V) highlights the benefits of prompt diversity within the self-generated framework. Utilizing diverse prompts (A1) rather than a single fixed prompt (A3) enriches data diversity, thereby enhancing generalization. Notably, these strong zero-shot generalization results are achieved without requiring any task-specific instruction pairs. Even when the training data is constructed using only randomly sampled prompts, the model successfully leverages the inherent capabilities of the LLM backbone to perform well. Furthermore, we explore the impact of model strength by using Llama3.1-70B to generate training targets (B3). Although the lower perplexity (2.20) suggests a closer data distribution, this setting does not consistently yield improvements across all tasks compared to the self-generated experiment (A1).
@@ VII-B With LoRA Adapter
In the LoRA adapter setting, we introduce trainable parameters to the backbone LLM, which is expected to increase model capacity and help mitigate the distribution mismatch problem. In the self-generation setup (C1), we find that adding LoRA layers yields similar or only slightly better performance. This indicates that, under self-generation settings, incorporating LoRA adapters does not provide significant advantages. In other words, fine-tuning a lightweight modality adapter is sufficient for cross-modal alignment. Interestingly, when training with Qwen2.5-generated data (C2), performances on audio processing tasks, such as Dynamic-SUPERB, MMAU, and SAKURA-Single, is comparable to the self-generation setup (A2). However, it experiences a significant degradation in SAKURA-Multi and Speech-IFEval, which requires additional text knowledge and instruction-following ability.
This disparity indicates that while LoRA adapters can mitigate distribution mismatch for in-domain tasks, they fail to preserve broader capabilities on out-of-domain benchmarks, suggesting that the primary utility of LoRA in such scenarios is to encode distributional mismatch between external targets and the backbone’s native behavior. This finding may provide a rationale for why previous works observed that adjusting or pruning LoRA weights during inference can help recover instruction-following ability [92, 72]. While introducing LoRA adapters remains a common approach in recent LALM development, our results suggest that reducing the discrepancy between training data and the model’s native distribution is a more critical factor for preserving generalization than architectural modifications alone.
@@ VIII Conclusion
In this work, we introduced DeSTA2.5-Audio, a general-purpose LALM built on the self-generated cross-modal alignment framework, DeSTA. By enabling the backbone LLM to generate its own training targets, we eliminate distributional shifts in the training dataset, thereby mitigating the catastrophic forgetting problem commonly observed in prior works. Under this framework, we successfully develop a robust and generalizable LALM without relying on task-specific instruction data. To support this, we constructed DeSTA-AQA5M, a large-scale, task-agnostic dataset comprising 5 million audio-text pairs from 50 diverse audio datasets. Trained solely on this corpus, DeSTA2.5-Audio achieves state-of-the-art or competitive results across benchmarks such as Dynamic-SUPERB (Phase-1 and Phase-2), MMAU, SAKURA, Speech-IFEval, and VoiceBench. Our analytical study reveals that the model effectively maps audio information into semantic concepts within the LLM representation space. This successful alignment enables the model to leverage its intrinsic instruction-following capabilities, thereby facilitating robust performance on downstream instruction-based benchmarks. Our findings introduce a new training philosophy for future LALM development, highlighting the importance of training data quality and consistency over quantity. Compared to previous studies, our self-generation strategy offers a scalable and robust solution with superior generalization.
While our current design leverages text descriptions as an intermediate bridge, we acknowledge that not all acoustic nuances can be effectively captured through textual representations. In future work, we will explore audio-to-audio supervision and preference learning from human feedback to better capture subtle, non-text-expressible acoustic features, advancing toward LALMs with more complete and human-like audio perception.